# Fault Diagnosis in New Wind Turbines using Knowledge from Existing Turbines by Generative Domain Adaptation

### v1

@@ 2.3 Domain mapping for unsupervised anomaly detection
Few studies exist on domain mapping applications for unsupervised anomaly detection, even beyond the wind energy research field. Hardly any domain adaptation studies exist for anomaly detection tasks as in the present study where only normal training samples are available combined with limited data in the target domain [yangAnomalyDetectionDomain2023]. To our knowledge, our study is the first to propose a CycleGAN-based approach to unsupervised anomaly detection using domain mapping.
Moreover, only few works have even performed the underlying domain translation with time series data, i.e., mapping time series, regardless of the task and outside the scope of wind turbine SCADA data. CycleGAN is proposed for data augmentation by generating artificial damaged states of acceleration data in [luleciCycleGANUndamagedtodamagedDomain2023a] through a mapping of undamaged-to-damaged conditions. Pattnaik et al. [pattnaikCycleGANBasedUnsupervised2022] translates time series across machines for bearing fault diagnosis using a CycleGAN framework. Unlike our anomaly detection study with scarce target data, their model maps abundant but unlabeled target domain data to the source domain where it is subsequently used in a fault classifier pretrained on the source domain. Results showed a significant outperformance compared to conventional domain adaptation methods when the domain shift is large, i.e., when mapping from substantially different but related machines.
@@ 3 Dataset
Our available raw data comprises 10-minute averaged SCADA data of 7 wind turbines, each from a distinct wind farm. While the turbines share the same manufacturer to ensure a matching SCADA variable system, they exhibit significantly different characteristics, model types, power rating, maintenance and fault history, and geographical location. An incident log is further provided for each WT. These logs contain a binary flag for each 10-minute measurement, describing time frames during which unknown and unspecified incidents occurred in the WT operation. While these incidents do not necessarily represent anomalous behavior or faults, we consider no measurements during these marked time frames as coming from normal operation behavior. A more detailed overview and an illustration of the 7 WTs is provided in Appendix D.
@@
Our study considers source-target pairs of WTs for domain adaptation, with one WT defining the source domain and the pair’s second WT defining the target domain. Our goal is to map SCADA samples between the source domain, consisting of a randomly selected WT, and the target domain, represented by another WT of a different WT type from a different wind farm. The data from the source domain WT is split into a training set (first 70% of the data), a validation set (30% of the training data), and a test set (the last 30% of data). For the target domain WT, starting with its full dataset, we first apply the same 70-30-30 split as with the source WT. As a next step we apply a data scarcity scenario, by artificially shortening the training and validation set. Specifically, we shorten the training/validation set to consecutive time frames of 1 to 8 weeks immediately preceding the test set, which remains the same across all scarcity scenarios. To ensure dataset size consistency when comparing across different WTs, we always refer to one week of data to comprise $1008$ (6 10-minute sample per hour * 24 * 7 days) SCADA samples instead of selecting by calendar week, as maintenance, data gaps, or incidents could cause strongly changing dataset sizes across turbines.
@@
These datasets are further normalized according to statistics calculated over the training set of the source domain such that each SCADA variable falls into a value range of $[-1,1]$. The target domain data is normalized using the same min-max normalization formula while using source domain statistics and therefore not confined within this range.
@@
From each set, we apply a sliding window approach to extract a sample dataset consisting of 12-hour samples of consecutive SCADA measurements with 11 variables. Selected were the mean, maximum, and minimum wind speed, rotor speed, power output values, as well as two temperatures from internal components, namely the mean stator and rotor temperature. Due to the one-to-one mapping inherent with CycleGAN [manytomanyGAN], we refrained from using variables with a larger stochastic component such as the ambient temperature. For instance, an idle WT operational state, represented through our highly correlated 11 input features, could exist with numerous possible ambient temperature variations, which may necessitate probabilistic domain mapping (many-to-many mappings, e.g. [manytomanyGAN]). The chosen 12-hours time frame allows us to better capture temporal dynamics within a sample, for instance thermal processes. An illustration of our dataset splits and resulting SCADA samples is shown in Figure 1.
@@
Normal behavior models require that only normal samples, i.e., only healthy WT operation states, are used for training. We apply WT-specific filtering to exclude possibly abnormal samples from the training and validation sets. First, we exclude all measurements that fall into a time frame marked in the incident logs. Moreover, we apply a rated power filter rule and Mahalanobis distance-based filtering based on [mckinnonComparisonNovelSCADA2022] to remove outliers, measurements from curtailment, and other outliers. No filtering is performed on the test sets.
@@ 4 Methodology
Our study aims to improve the reliability of fault diagnosis models for WTs with unrepresentative training data by leveraging models and data from similar, but not identical, WTs for which representative training data is available. To achieve this, we employ a generative domain adaptation approach, namely domain mapping. This section outlines our machine learning models, domain mapping losses, and evaluation metrics used. Our proposed framework consists of a domain mapping network that can translate source domain samples (consecutive 12-hour SCADA samples) into the target domain and vice versa. Despite the translation being learned cyclically and in both directions during training, we are ultimately interested in mapping (scarce) target domain data to the source domain for subsequent anomaly detection. The anomaly detection underlying the fault diagnosis of the mapped target domain data is finally achieved with an autoencoder-based NBM pretrained on the source domain. The proposed workflow is shown in Figure 2.
@@ 4.1 Autoencoder-based NBM
We employ an NBM based on autoencoders [liComprehensiveSurveyDesign2023] for anomaly detection and fault diagnosis. Autoencoders are unsupervised neural networks trained to encode input data into a lower dimensional representation (encoder) and to subsequently reconstruct the input data from the compressed representation (decoder). To reconstruct the input as exactly as possible, an autoencoder learns to encode the most important features, thereby learning critical variable and feature relationships. An input sample that significantly deviates from the training set (i.e., an anomaly) will thus result at test time in an elevated reconstruction error, representing an anomaly score. Autoencoders have been employed as WT-specific NBMs (e.g., [jonasVibrationFaultDetection2023, roelofsAutoencoderbasedAnomalyRoot2021, renstromSystemwideAnomalyDetection2020]). Our autoencoder takes as input 12-hour SCADA samples with 11 channels (792 data points) and learns to reconstruct them based on an encoding size (bottleneck dimension) of 72 units. A model is trained on only normal source domain training samples for each WT domain pair. At test time, we categorize all samples with a reconstruction error (mean absolute error between the autoencoder reconstruction output and its original input) above the defined WT-specific threshold as anomalous. The threshold is set to detect far out outliers [schwertmanSimpleMoreGeneral2004] for all models as $T=q_{3}+3\left(q_{3}-q_{1}\right)$, where $q_{3}$ and $q_{1}$ represent the 75th and 25th percentile of the normal validation data reconstruction errors, respectively. Further information about the autoencoder model is outlined in Appendix A.
@@ 4.2 Domain mapping model
Our domain mapping network is based on the CycleGAN formulation [zhuUnpairedImagetoImageTranslation2017]. The network consists of two one-directional generative adversarial networks (section 4.3) taking as input SCADA samples of one WT to transform them into data resembling the other WT. The generators are trained to synthesize realistic samples in an adversarial way by trying to outperform the WT-specific critics, which try to distinguish between fake (generated) and real samples of their particular WT. Unconstrained generators, as in the case of [jinConditionMonitoringWind2023], can however potentially transform an input SCADA sample in any way and disregard the inherent content within the sample, which represents a particular state of the WT, e.g., idle power, maximal power production, or anomalous behavior. We propose and demonstrate that adding consistency losses (section 4.4) to the network enforces the content-preservation for anomaly detection. Most importantly, a cycle-consistency loss enforces that a sample first mapped to the other domain and then back to its original domain should remain similar. Figure 3 illustrates the concept of our domain mapping network.
@@ 4.3 Generators and critics
The framework consists of two generators $G_{ST},G_{TS}$ and two discriminators ${Disc}_{T},{Disc}_{S}$. The generators are tasked to map a sample from one domain, which is their input, to resemble a sample from the other domain. There is one generator for each domain direction, i.e., $G_{ST}:\mathcal{D}_{\mathcal{S}}\rightarrow\mathcal{D}_{\mathcal{T}},G_{TS}:\mathcal{D}_{\mathcal{T}}\rightarrow\mathcal{D}_{\mathcal{S}}$. The generator $G_{ST}$ maps an input SCADA sample from $\mathcal{D}_{\mathcal{S}}$ to resemble a sample belonging to domain $\mathcal{D}_{\mathcal{T}}$. As opposed to the standard GAN formulation, the generators take as inputs real samples of a domain to translate, instead of synthesizing samples from random noise. The discriminators generally act as critics, assessing whether a sample belongs to the underlying probability distribution or not. ${Disc}_{T}\left(x\right)$ is generally a score reflecting whether a sample candidate $x$ belongs to the corresponding domain distribution $P_{T}$ or not. In our work, we use the GAN-QP framework [suGANQPNovelGAN2018, liSystematicSurveyRegularization2023] as in [jinConditionMonitoringWind2023] to train the generators and critics to generate realistic samples. Further details about model architectures, hyperparameter optimization, and complete training procedure are presented in Appendix B.
@@ 4.4 Content preservation losses
As the goal is to ultimately perform fault diagnosis with mapped data, it is critical to ensure that the generator mapping preserves the SCADA sample content, i.e., the operational state represented by the sample such as constant maximum power output and in particular anomalous behavior. The proposed cycle-consistency loss in CycleGAN [zhuUnpairedImagetoImageTranslation2017] constrains the domain mapping to encourage that a mapped sample mapped back to its original domain resembles the original sample. Formally, it encourages $G_{ST}\left(G_{TS}\left(t\right)\right)\approx t$ and vice versa. This is achieved by adding an L1 loss to the generator loss punishing deviations between original and cycled samples:
@@
However, the cycle-consistency loss can be in our case insufficient to ensure a consistent content mapping across domains. To illustrate this, let us consider adding a hypothetical flipping operation to the generators. Typically, with image data, flipping an input will cause discriminators to reject the mapping. For instance, flipping a horse image upside down before translating it into a zebra will never generate a realistic zebra, whereas “flipping” a SCADA input (e.g., maximum power mapped to zero power) before translation can still generate realistic domain samples. As the same flipping operation is performed in the opposite domain direction, the cycled output will resemble the original input, thereby still enforcing the cycle-consistency but without actually preserving content. To account for this, we further restrict the mapping space with additional consistency losses.
@@ 4.5 Evaluation and benchmarks
[…]
@@ 5 Results and Discussion
5.1 Fault diagnosis performance
We assess the fault diagnosis performance of our proposed domain mapping network in terms of producing similar anomaly scores as a target WT NBM trained on abundant data, as outlined in Figure 5. In total, we evaluated 28 source-target WT domain pairs using 6 distinct target domain WTs, as one of the seven WTs (WT07) was used during model selection and hyper-parameter optimization (see Appendix B), with varying degrees of target training data scarcity of 1-8 weeks. All domain pairs consist of a source and target WT differing in turbine specifications in terms of model and having a unique rated power capacity (see Appendix D), farm, and geographical location. For each source-target domain pair, we compared the performance to an NBM trained on scarce data only and a fine-tuning benchmark.
@@ 5.1 Fault diagnosis performance
For all 6 target WTs considered in our study, we first trained NBMs using only scarce target training data under varying scarcity scenarios. That is, we trained an NBM using 1-8 weeks of training data and evaluated the resulting anomaly scores on the fixed test set in comparison to the ground truth. Our findings in Table 1 show that NBMs trained on limited training sets achieved comparably poor fault diagnosis performance, with the mean F1-score generally decreasing as less training data becomes available. While a training set comprising 2 months of data achieved on average an F1-score of 84.3% across all 6 target WTs, the mean score substantially decreased to 51.3% when only 2 weeks are considered, showing a strong drop in similarity of anomaly scores compared to the NBM trained on a full training dataset. Limited data can lead to unrepresentative training sets with an insufficient coverage of samples from normal operational behavior. An example of representative and unrepresentative training data is shown in Appendix E. Thus, the autoencoder-based NBM becomes incapable of reconstructing normal data at test time that was missing from the training set. Our results indicating that training NBMs under data scarcity can result in unreliable fault diagnosis support the findings of e.g., [grataloupWindTurbineCondition2024, jenkelPrivacyPreservingFleetWideLearning2023].
Improving fault diagnosis with domain adaptation.
To overcome these limitations, we employed domain adaptation techniques. We trained and evaluated our proposed domain mapping technique and fine-tuning as comparison benchmark to overcome these limitations. We present the achieved improvements in F1-score compared to training an NBM on scarce data only (i.e., baseline without domain adaptation) for each domain pair in Table 2, further averaged by data scarcity degree. The absolute F1-scores are available in Table A4 in Appendix C. Both fine-tuning and our domain mapping network achieved an improvement in fault diagnosis performance in scenarios with strong data scarcity, i.e., particularly in situations where the scarce NBM fails to achieve reliable fault diagnosis.
@@
As shown in Table 2, both domain adaptation methods managed to to improve fault diagnosis performance over training on scarce data alone in most domain pairs and across almost all data scarcity degrees. Fine-tuning was able to improve the performance by increasing the F1-score on average by +4.7% points when only 1 month of target training data was available and +9.3% with 2 weeks of data, showing improvements across all scarcity scenarios except 2 months. These results further suggest that fine-tuning can be employed to adapt knowledge embedded in the source domain NBM to a limited target domain.
Notably, our proposed domain mapping approach outperformed fine-tuning across all considered training set sizes, especially in scenarios with severe data scarcity. In the scenarios ranging from 1 week to 1 month of data, our domain mapping approach achieved a substantial performance gain over not only the NBM trained on scarce data but also consistently over fine-tuning, with an increase in mean F1 score by +10.3% points for 1 month and +16.8% for 2 weeks. Our presented domain mapping approach makes use of the entire source WT’s dataset to learn mappings and uses its reliable model for anomaly detection, as opposed to fine-tuning, which exclusively relies on only the trained source NBM parameters being adjusted by a limited target training set, thereby limiting its acquired knowledge of the domain shift. Our findings highlight the effectiveness and potential of our domain mapping method as an alternative approach to conventional fine-tuning to mitigate the challenges of data scarcity in WT fault diagnosis. Our proposed technique can enable more reliable and earlier fault diagnosis, for instance for newly installed wind turbines, by incorporating models and data of only one reference source wind turbine with abundant data.
@@
Lastly, we note strong differences in performance across target WTs. These could be attributed to the randomness of the selected training set weeks (e.g., contained weather conditions, variety of operational states), the test set characteristics (e.g., number of faults), general turbine behavior, and the domain pair, exhibiting varying domain shifts across different WTs. Further research is needed to investigate these variations and to possibly adjust the network and hyperparameters to the specific characteristics of WTs. For certain domain pairs, fine-tuning may retrospectively turn out to be a better solution compared to domain mapping. In some few other cases, both domain adaptation methods may even lead to a substantial drop in performance even with limited training data. Further investigations are required to clearly identify these causes and for finding mitigation strategies. These variations may cause uncertainties regarding an appropriate method selection, as due to the fundamental lack of knowledge about the test set, it remains unclear which approach to choose at training time.

### v2

@@ 2.3 Domain mapping for unsupervised anomaly detection
Few studies exist on domain mapping applications for unsupervised anomaly detection, even beyond the wind energy research field. Hardly any domain adaptation studies exist for anomaly detection tasks as in the present study where only normal training samples are available combined with limited data in the target domain [15]. To our knowledge, our study is the first to propose a CycleGAN-based approach to unsupervised anomaly detection for WT fault detection using domain mapping.
Moreover, only few works have even performed the underlying domain translation with time series data, i.e., mapping time series, regardless of the task and outside the scope of wind turbine SCADA data. CycleGAN is proposed for data augmentation by generating artificial damaged states of acceleration data in [33] through a mapping of undamaged-to-damaged conditions. Pattnaik et al. [20] translates time series across machines for bearing fault diagnosis using a CycleGAN framework. Unlike our anomaly detection study with scarce target data, their model maps abundant but unlabeled target domain data to the source domain where it is subsequently used in a fault classifier pretrained on the source domain. Results showed a significant outperformance compared to conventional domain adaptation methods when the domain shift is large, i.e., when mapping from substantially different but related machines.
@@ 3 Dataset
Our available raw data comprises 10-minute averaged SCADA data of 7 wind turbines, each from a distinct wind farm. While the turbines share the same manufacturer to ensure a matching SCADA variable system, they exhibit significantly different characteristics, model types, power rating, maintenance and fault history, and geographical location. An incident log is further provided for each WT. These logs contain a binary flag for each 10-minute measurement, describing time frames during which unknown and unspecified incidents occurred in the WT operation. These incidents do not necessarily represent anomalous behavior or faults. Thus, we do not use incidents for evaluation or classification but only for additional filtering, as we consider no measurements during these marked time frames as coming from normal operation. A more detailed overview and an illustration of the 7 WTs is provided in Appendix D.
@@
Our study considers source-target pairs of WTs for domain adaptation, with one data-rich WT defining the source domain and the pair’s second data-scarce WT defining the target domain. Our goal is to map SCADA samples between the source domain, consisting of a randomly selected WT, and the target domain, represented by another WT of a different WT type from a different wind farm. The data from the source domain WT is split into a training set (first 70% of the data), a validation set (30% of the training data), and a test set (the last 30% of data). For the target domain WT, starting with its full dataset, we first apply the same 70-30-30 split as with the source WT. As a next step we apply a data scarcity scenario, by artificially shortening the training and validation set. Specifically, we shorten the training/validation set to consecutive time frames of 1 to 8 weeks immediately preceding the test set, which remains the same across all scarcity scenarios. To ensure dataset size consistency when comparing across different WTs, we always refer to one week of data to comprise $1008$ (6 10-minute sample per hour * 24 * 7 days) SCADA samples instead of selecting by calendar week, as maintenance, data gaps, or incidents could cause strongly changing dataset sizes across turbines.
@@
These datasets are further normalized according to statistics calculated over the training set of the source domain such that each SCADA variable falls into a value range of [1,1][-1,1]. The target domain data is normalized using the same min-max normalization formula but using source domain statistics and therefore not confined within this range.
@@
From each set, we apply a sliding window approach to extract a sample dataset consisting of 12-hour samples of consecutive SCADA measurements with 11 variables. Selected were the mean, maximum, and minimum wind speed, rotor speed, power output values, as well as two temperatures from internal components, namely the mean stator and rotor temperature. Due to the one-to-one mapping inherent with CycleGAN [34], we refrained from using variables with a larger stochastic component such as the ambient temperature. For instance, an idle WT operational state, represented through our highly correlated 11 input features, could exist with numerous possible ambient temperature variations, which may necessitate probabilistic domain mapping (many-to-many mappings, e.g. [34]). The chosen 12-hours time frame allows us to better capture temporal dynamics within a sample, for instance thermal processes. An illustration of our dataset splits and resulting SCADA samples is shown in Figure 1.
@@
Normal behavior models require that only normal samples, i.e., only healthy WT operation states, are used for training. We apply WT-specific filtering to exclude possibly abnormal samples from the training and validation sets. First, we exclude all measurements that fall into a time frame marked in the incident logs. Moreover, we apply a rated power filter rule and Mahalanobis distance-based filtering based on [35] to remove outliers and measurements from curtailment. No filtering is performed on the test sets.
@@ 4 Methodology
Our study aims to improve the reliability of fault detection models for WTs with unrepresentative training data by leveraging models and data from similar, but not identical, WTs for which representative training data is available. To achieve this, we employ a generative domain adaptation approach, namely domain mapping. This section outlines our machine learning models, domain mapping losses, and evaluation metrics used. Our proposed framework consists of a domain mapping network that can translate source domain samples (consecutive 12-hour SCADA samples) into the target domain and vice versa. Despite the translation being learned cyclically and in both directions during training, we are ultimately interested in mapping data from the data-scarce target WT to the data-rich source WT for subsequent anomaly detection. Once the data is mapped, the anomaly detection is achieved through evaluations with an autoencoder-based NBM pretrained on the source domain. The proposed workflow is shown in Figure 2.
@@ 4.1 Autoencoder-based NBM
We employ an NBM based on autoencoders [36] for anomaly detection and fault detection. Autoencoders are unsupervised neural networks trained to encode input data into a lower dimensional representation (encoder) and to subsequently reconstruct the input data from the compressed representation (decoder). To reconstruct the input as exactly as possible, an autoencoder learns to encode the most important features, thereby learning critical variable and feature relationships. An input sample that significantly deviates from the training set (i.e., an anomaly) will thus result at test time in an elevated reconstruction error, representing an anomaly score. Autoencoders have been employed as WT-specific NBMs (e.g., [37, 38, 39]). Our autoencoder takes as input 12-hour SCADA samples with 11 channels (792 data points) and learns to reconstruct them based on an encoding size (bottleneck dimension) of 72 units. A model is trained on only normal source domain training samples for each WT domain pair. At test time, we categorize all samples with a reconstruction error (mean absolute error between the autoencoder reconstruction output and its original input) above the defined WT-specific threshold as anomalous. The threshold is set to detect far out outliers [40] for all models as $T=q_{3}+3\left(q_{3}-q_{1}\right)$, where $q_{3}$ and $q_{1}$ represent the 75th and 25th percentile of the normal validation data reconstruction errors, respectively. Further information about the autoencoder model is outlined in Appendix A.
@@ 4.2 Domain mapping model
Our domain mapping network is based on the CycleGAN formulation [21]. The network consists of two one-directional generative adversarial networks (section 4.3) taking as input SCADA samples of one WT to transform them into data resembling the other WT. The generators are trained to synthesize realistic samples in an adversarial way by trying to outperform the WT-specific critics, which try to distinguish between fake (generated) and real samples of their particular WT. Unconstrained generators, as in the case of [19], can however potentially transform an input SCADA sample in any way and disregard the inherent content within the sample, which represents a particular state of the WT, e.g., idle power, maximal power production, or anomalous behavior. We propose and demonstrate that adding consistency losses (section 4.4) to the network enforces the content-preservation for anomaly detection. Most importantly, a cycle-consistency loss enforces that a sample first mapped to the other domain and then back to its original domain should remain similar. Figure 3 illustrates the concept of our domain mapping network.
@@ 4.3 Generators and critics
The framework consists of two generators $G_{ST},G_{TS}$ and two discriminators ${Disc}_{T},{Disc}_{S}$. The generators are tasked to map a sample from one domain, which is their input, to resemble a sample from the other domain. There is one generator for each domain direction, i.e., $G_{ST}:\mathcal{D}_{\mathcal{S}}\rightarrow\mathcal{D}_{\mathcal{T}},G_{TS}:\mathcal{D}_{\mathcal{T}}\rightarrow\mathcal{D}_{\mathcal{S}}$. The generator $G_{ST}$ maps an input SCADA sample from $\mathcal{D}_{\mathcal{S}}$ to resemble a sample belonging to domain $\mathcal{D}_{\mathcal{T}}$. As opposed to the standard GAN formulation, the generators take as inputs real samples of a domain to translate, instead of synthesizing samples from random noise. The discriminators generally act as critics, assessing whether a sample belongs to the underlying probability distribution or not. ${Disc}_{T}\left(x\right)$ is generally a score reflecting whether a sample candidate $x$ belongs to the corresponding domain distribution $P_{T}$ or not. In our work, we use the GAN-QP framework [41, 42] as in [19] to train the generators and critics to generate realistic samples. Further details about model architectures, hyperparameter optimization, and complete training procedure are presented in Appendix B.
@@ 4.4 Content preservation losses
As the goal is to ultimately perform fault detection with mapped data, it is critical to ensure that the generator mapping preserves the SCADA sample content, i.e., the operational state represented by the sample such as constant maximum power output and in particular anomalous behavior. The proposed cycle-consistency loss in CycleGAN [21] constrains the domain mapping to encourage that a mapped sample mapped back to its original domain resembles the original sample. This constraint aims to ensure that during a cycled mapping there were no changes or features added or removed. As an example, a zebra mapped to a horse and then back should remain the same animal in the same position and field, as otherwise content alterations may have been performed by the models. Formally, it encourages $G_{ST}\left(G_{TS}\left(t\right)\right)\approx t$ and vice versa. This is achieved by adding an L1 loss to the generator loss punishing deviations between original and cycled samples:
@@
However, the cycle-consistency loss can be in our case insufficient to ensure a consistent content mapping across domains. To illustrate this, let us consider adding a hypothetical flipping operation to the generators. Typically, with image data, flipping an input will cause discriminators to reject the mapping. For instance, flipping a horse image upside down before translating it into a zebra will never generate a realistic zebra, whereas “flipping” a SCADA input (e.g., maximum power mapped to zero power) before translation can still generate realistic domain samples. If the same flipping operation is performed in the opposite domain direction, the cycled output will resemble the original input, thereby still enforcing the cycle-consistency but without actually preserving content. To account for this, we further restrict the mapping space with additional consistency losses.
@@ 4.5 Evaluation and benchmarks
$\text{F1-Score}=\frac{2\text{TP}}{2\text{TP}+\text{FP}+\text{FN}}$ (4)
@@ 5 Results and Discussion
5.1 Fault detection performance
We assess the fault detection performance of our proposed domain mapping network in terms of producing similar anomaly scores as a target WT NBM trained on abundant data, as outlined in Figure 5. In total, we evaluated 28 source-target WT domain pairs using 6 distinct target domain WTs, as one of the seven WTs (WT07) was used during model selection and hyper-parameter optimization (see Appendix B), with varying degrees of target training data scarcity of 1-8 weeks. All domain pairs consist of a source and target WT differing in turbine specifications in terms of model and having a unique rated power capacity (see Appendix D), farm, and geographical location. For each source-target domain pair, we compared the performance to an NBM trained on scarce data only and a fine-tuning benchmark.
@@ 5.1 Fault detection performance
For all 6 target WTs considered in our study, we first trained NBMs using only scarce target training data under varying scarcity scenarios. That is, we trained an NBM using 1-8 weeks of training data and evaluated the resulting anomaly scores on the fixed test set in comparison to the ground truth. Our findings in Table 1 show that NBMs trained on limited training sets achieved comparably poor fault detection performance, with the mean F1-score generally decreasing as less training data becomes available. While a training set comprising 2 months of data achieved on average an F1-score of 84.3% across all 6 target WTs, the mean score substantially decreased to 51.3% when only 2 weeks are considered, showing a strong drop in similarity of anomaly scores compared to the NBM trained on a full training dataset. Limited data can lead to unrepresentative training sets with an insufficient coverage of samples from normal operational behavior. An example of representative and unrepresentative training data is shown in Appendix E. Thus, the autoencoder-based NBM becomes incapable of reconstructing normal data at test time that was missing from the training set. Our results indicating that training NBMs under data scarcity can result in unreliable fault detection support the findings of e.g., [7, 8].
Improving fault detection with domain adaptation.
To overcome these limitations, we employed domain adaptation techniques. We trained and evaluated our proposed domain mapping technique and fine-tuning as comparison benchmark to overcome these limitations. We present the achieved improvements in F1-score compared to training an NBM on scarce data only (i.e., baseline without domain adaptation) for each domain pair in Table 2, further averaged by data scarcity degree. The absolute F1-scores are available in Table A4 in Appendix C. Both fine-tuning and our domain mapping network achieved an improvement in fault detection performance in scenarios with strong data scarcity, i.e., particularly in situations where the scarce NBM fails to achieve reliable fault detection.
@@
As shown in Table 2, both domain adaptation methods managed to to improve fault detection performance over training on scarce data alone in most domain pairs and across almost all data scarcity degrees. Fine-tuning was able to improve the performance by increasing the F1-score on average by +4.7% points when only 1 month of target training data was available and +9.3% with 2 weeks of data, showing improvements across all scarcity scenarios except 2 months. These results further suggest that fine-tuning can be employed to adapt knowledge embedded in the source domain NBM to a limited target domain.
Notably, our proposed domain mapping approach outperformed fine-tuning across all considered training set sizes, especially in scenarios with severe data scarcity. In the scenarios ranging from 1 week to 1 month of data, our domain mapping approach achieved a substantial performance gain over not only the NBM trained on scarce data but also consistently over fine-tuning, with an increase in mean F1 score by +10.3% points for 1 month and +16.8% for 2 weeks. Our presented domain mapping approach makes use of the entire source WT’s dataset to learn mappings and uses its reliable model for anomaly detection, as opposed to fine-tuning, which exclusively relies on only the trained source NBM parameters being adjusted by a limited target training set, thereby limiting its acquired knowledge of the domain shift. Our findings highlight the effectiveness and potential of our domain mapping method as an alternative approach to conventional fine-tuning to mitigate the challenges of data scarcity in WT fault detection. Our proposed technique can enable more reliable and earlier fault detection, for instance for newly installed wind turbines, by incorporating models and data of only one reference source wind turbine with abundant data.
@@
Lastly, we note strong differences in performance across target WTs. These could be attributed to the randomness of the selected training set weeks (e.g., weather conditions, variety of operational states), test set characteristics (e.g., number of faults), or general turbine behavior. Further research is needed to investigate these variations and to possibly adjust the network and hyperparameters to the specific characteristics of WTs. Furthermore, our findings show significant variations across WT pairs. A possible cause of performance differences may be, among other, that WT pairs can exhibit varying domain shifts. Thus, selecting suitable source-target pairs can be important. However, the fundamental unavailability of representative validation data impedes an informed pair selection. An exploration is needed to investigate and assess the suitability of WT pairs for domain mapping, which may involve e.g., a quantification of the domain shift. For certain domain pairs, fine-tuning may also retrospectively turn out to be a better solution compared to domain mapping. In some few other cases, both domain adaptation methods may even lead to a substantial drop in performance even with limited training data. Further investigations are required to clearly identify these causes and for finding mitigation strategies. These variations may cause uncertainties regarding an appropriate method selection, as due to the inherent lack of knowledge about the test set, it remains unclear which approach to choose at training time.